OpenStudy (anonymous):

what is the square root of 81/49?

hero (hero):

\[\sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\]

hero (hero):

So: \[\sqrt{\frac{81}{49}} = \frac{\sqrt{81}}{\sqrt{49}}\]

OpenStudy (anonymous):

9/7